# Dynamics of the Morse vector field

### v1

@@ Abstract.
The (negative) gradient vector fields of Morse functions on a compact manifold provide an important example in dynamical system. In this note we prove two important properties of this kind of vector field: Connectedness of critical points through orbits and exponential shrinkage of the flow on stable manifolds. We also find applications in showing some vanishing results of maps or curvature operators.
@@ 1. Introduction
,which are diffeomorphic to Euclidean open balls of dimension $\mathrm{ind}_{f}(p)$ and $\mathrm{dim}(M)-\mathrm{ind}_{f}(p)$.
@@ 2. Main results
In order to take covariant derivatives of the cross section $\mathrm{d}\phi_{t}(v)=\mathrm{d}\Phi|_{(\cdot~,t)}(v,0)$, we introduce the pull-back bundle $\Phi^{-1}(TM)$ of $TM$ and the pull-back connection $\tilde{\nabla}$ of Levi-Civita connection $\nabla$ on $TM$. The bundle $\Phi^{-1}(TM)\to M$ admits a natural definition for which there exists a $C^{\infty}$-linear morphism $\psi:\Phi^{-1}(TM)\to TM$ such that the diagram
@@
$(\tilde{\nabla}\mathrm{d}\Phi)(X,Y)=\tilde{\nabla}_{X}\mathrm{d}\Phi(Y)-\tilde{\nabla}_{Y}\mathrm{d}\Phi(X)-\mathrm{d}\Phi([X,Y])=0.$
@@
Notice that $\mathrm{d}\phi_{t}|_{x}(v_{x})\in T_{\phi_{t}(x)}W^{s}(p)$ since $v_{x}\in T_{x}W^{s}(p)$ - a consequence of the invariance of $W^{s}(p)$ under the flow $\phi_{t}$, so it converges to a vector in $T_{p}W^{s}(p)$. Moreover $\text{Hess}(f)|_{p}$ is positive definite on $T_{p}W^{s}(p)$ , i.e. $\exists~c>0~s.t.~\text{Hess}(f)|_{p}(w,w)\geq c|w|^{2}$ for $\forall w\in T_{p}W^{s}(p)$. Thus for suffciently large $t$, $\text{Hess}(f)|_{\phi_{t}(x)}(\mathrm{d}\phi_{t}|_{x}(v_{x}),\mathrm{d}\phi_{t}|_{x}(v_{x}))\geq\frac{c}{2}|\mathrm{d}\phi_{t}|_{x}(v_{x})|^{2}=cE(t)$. It’s basically a linear-algebraic fact: If a quadratic form $B$ is positive definite on a subspace $V$, then small perturbations of $B$ and $V$ still preserve the positivity.
@@ 3. Applications
for $\forall~Y_{1},\cdots,Y_{r}\in\Gamma(M,TM)$. A direct computation analogous to the case of normal Lie derivative gives
$\mathcal{L}_{X}^{D}\omega(Y_{1},\cdots,Y_{r})=D_{X}\big(\omega(Y_{1},\cdots,Y_{r})\big)-\displaystyle\sum_{i=1}^{r}\omega(Y_{1},\cdots,[X,Y_{i}],\cdots,Y_{r}),$
and we obtain a similar Cartan formula: $\mathcal{L}^{D}_{X}=\imath_{X}d_{D}+d_{D}\imath_{X}$, where $d_{D}:\Omega^{\bullet}(M,E)\to\Omega^{\bullet+1}(M,E)$ is the induced exterior derivative of the connection $D$. An explicit formula for $d_{D}$ can be found in [4].
One final point to mention is the endomorphism bundle $\text{End}(E)$, together with its connection $\tilde{D}$ induced by the connection $D$ on $E$. Note that the curvature operator $R^{D}\in\Omega^{2}(M,\text{End}(E))$ and we have the well-known Bianchi identity $d_{\tilde{D}}R^{D}=0$.
@@
Remark: Notice that the condition $\mathcal{L}^{\tilde{D}}_{\nabla f}R^{D}=d_{\tilde{D}}\imath_{\nabla f}R^{D}=0$ is weaker than the condition $\imath_{\nabla f}R^{D}=0$ in Proposition 3.4 of [4], owing to the Cartan formula and Bianchi identity above. The proof is also different.

### v2

@@ Abstract.
The (negative) gradient vector fields of Morse functions on a compact manifold provide an important example in dynamical system. In this note we prove two important properties of this kind of vector field: Connectedness of critical points through orbits and exponential shrinkage of the flow on stable submanifolds. We also find applications in showing some vanishing results for maps or curvature operators.
@@ 1. Introduction
,which are diffeomorphic to Euclidean spaces of dimension $\mathrm{ind}_{f}(p)$ and $\mathrm{dim}(M)-\mathrm{ind}_{f}(p)$.
@@ 2. Main results
In order to take covariant derivatives of the cross section $\mathrm{d}\phi_{t}(v)=\mathrm{d}\Phi|_{(\cdot~,t)}(v,0)$, we introduce the pull-back bundle $\Phi^{-1}(TM)$ of $TM$ and the pull-back connection $\tilde{\nabla}$ of the Levi-Civita connection $\nabla$ on $TM$. The bundle $\Phi^{-1}(TM)\to M$ admits a natural definition for which there exists a $C^{\infty}$-linear morphism $\psi:\Phi^{-1}(TM)\to TM$ such that the diagram
@@
$(\mathrm{d}_{\tilde{\nabla}}\mathrm{d}\Phi)(X,Y)=\tilde{\nabla}_{X}\mathrm{d}\Phi(Y)-\tilde{\nabla}_{Y}\mathrm{d}\Phi(X)-\mathrm{d}\Phi([X,Y])=0.$
@@
Notice that $\mathrm{d}\phi_{t}|_{x}(v_{x})\in T_{\phi_{t}(x)}W^{s}(p)$ since $v_{x}\in T_{x}W^{s}(p)$ - a consequence of the invariance of $W^{s}(p)$ under the flow $\phi_{t}$, and $T_{\phi_{t}(x)}W^{s}(p)$ converges to $T_{p}W^{s}(p)$ as $t\to+\infty$. Moreover $\text{Hess}(f)|_{p}$ is positive definite on $T_{p}W^{s}(p)$ , i.e. $\exists~c>0~s.t.~\text{Hess}(f)|_{p}(w,w)\geq c|w|^{2}$ for $\forall w\in T_{p}W^{s}(p)$. Thus for suffciently large $t$, $\text{Hess}(f)|_{\phi_{t}(x)}(\mathrm{d}\phi_{t}|_{x}(v_{x}),\mathrm{d}\phi_{t}|_{x}(v_{x}))\geq\frac{c}{2}|\mathrm{d}\phi_{t}|_{x}(v_{x})|^{2}=cE(t)$. It’s basically a linear-algebraic fact: If a quadratic form $B$ is positive definite on a subspace $V$, then small perturbations of $B$ and $V$ still preserve the positivity.
@@ 3. Applications
for $\forall~Y_{1},\cdots,Y_{r}\in\Gamma(M,TM)$. It has a formula analogous to the normal Lie derivative:
$\mathcal{L}_{X}^{D}\omega(Y_{1},\cdots,Y_{r})=D_{X}\big(\omega(Y_{1},\cdots,Y_{r})\big)-\displaystyle\sum_{i=1}^{r}\omega(Y_{1},\cdots,[X,Y_{i}],\cdots,Y_{r})$
Proof.
It’s just an argument of Leibniz rule :
$\mathcal{L}_{X}^{D}\omega(Y_{1},\cdots,Y_{r})(x)=\lim_{t\to 0}\frac{1}{t}\Big(\tilde{\phi}_{-t}\big(\omega_{\phi_{t}(x)}\big(\mathrm{d}\phi_{t}(Y_{1}\big|_{x}),\cdots,\mathrm{d}\phi_{t}(Y_{r}\big|_{x})\big)\big)-\omega_{x}\big(Y_{1}\big|_{x},\cdots,Y_{r}\big|_{x})\Big)$
$=\lim_{t\to 0}\frac{1}{t}\Big(\tilde{\phi}_{-t}\big(\omega_{\phi_{t}(x)}(Y_{1}\big|_{\phi_{t}(x)},\cdots,Y_{r}\big|_{\phi_{t}(x)})\big)-\omega_{x}\big(Y_{1}\big|_{x},\cdots,Y_{r}\big|_{x}\big)\Big)+$
$\sum_{i=1}^{r}\lim_{t\to 0}\tilde{\phi}_{-t}\Big(\omega_{\phi_{t}(x)}\big(Y_{1}\big|_{\phi_{t}(x)},\cdots,\mathrm{d}\phi_{t}(\frac{Y_{i}|_{x}-\mathrm{d}\phi_{-t}(Y_{i}|_{\phi_{t}(x)})}{t}),\cdots,\mathrm{d}\phi_{t}(Y_{r}\big|_{x})\big)\Big)$
$=D_{X}\big(\omega(Y_{1},\cdots,Y_{r})\big)-\sum_{i=1}^{r}\omega\big(Y_{1},\cdots,\mathcal{L}_{X}(Y_{i}),\cdots,Y_{r}\big).$
Notice that it induces the normal formula by considering the trivial connection on the trivial line bundle over $M$. $\square$
Recall the induced exterior derivative $d_{D}:\Omega^{\bullet}(M,E)\to\Omega^{\bullet+1}(M,E)$ of the connection $D$ (see [4] for more details):
$d_{D}\omega\big(Y_{1},\cdots,Y_{r}\big)=\sum_{i=1}^{r}(-1)^{i-1}D_{Y_{i}}\big(\omega(Y_{1},\cdots,\hat{Y_{i}},\cdots,Y_{r})\big)$
$+\sum_{1\leq i,j\leq r}(-1)^{i+j}\omega\big([Y_{i},Y_{j}],Y_{1},\cdots,\hat{Y_{i}},\cdots,\hat{Y_{j}},\cdots,Y_{r}\big),$
we obtain a similar Cartan formula: $\mathcal{L}^{D}_{X}=\imath_{X}d_{D}+d_{D}\imath_{X}$, which is sufficient to verify for 0- and $1$-forms, since $\mathcal{L}_{X}^{D}$ is a derivation (of degree 0):
$(\imath_{X}d_{D}+d_{D}\imath_{X})f=d_{D}f(X)=D_{X}f=\mathcal{L}_{X}^{D}f;$
$(\imath_{X}d_{D}+d_{D}\imath_{X})\omega(Y)=d_{D}\omega(X,Y)+\big(d_{D}\omega(X)\big)Y$
$=D_{X}\omega(Y)-D_{Y}\omega(X)-\omega([X,Y])+D_{Y}\omega(X)=\mathcal{L}_{X}^{D}\omega(Y).$
One final point to mention is the endomorphism bundle $\text{End}(E)$, together with its connection $\tilde{D}=[d_{D},\cdot]$ induced by the connection $D$ on $E$. Note that the curvature operator $R^{D}\in\Omega^{2}(M,\text{End}(E))$ and we have the well-known Bianchi identity $d_{\tilde{D}}R^{D}=[d_{D},d_{D}^{2}]=0$.
@@
We conclude this note with two remarks. First, the condition $\mathcal{L}^{\tilde{D}}_{\nabla f}R^{D}=d_{\tilde{D}}\imath_{\nabla f}R^{D}=0$ is weaker than the condition $\imath_{\nabla f}R^{D}=0$ in Proposition 3.4 of [4], owing to the Cartan formula and Bianchi identity above, and the proof is also different. Second, by viewing $\mathrm{d}u$ as a $u^{-1}TN$ 1-form $\in\Omega^{1}(M,u^{-1}TN)$, we could get the same vanishing result in Proposition 3.1 with weaker condition:
Proposition 3.3.
If for a smooth map $u:M\to N$ and for some Morse function $f$ on $M$, $\mathcal{L}_{\nabla f}^{\tilde{\nabla}}du=0$, then $u$ must be a constant map. Here $\tilde{\nabla}$ is the pull-back connection of the Levi-Civita connection on $N$.
The proof is identical to that of Proposition 3.2, and the condition is indeed weaker for the same reason: Cartan formula and $\mathrm{d}_{\tilde{\nabla}}\mathrm{d}u=0$, which appears in the proof of Theorem 2.2.